# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v2

@@ Abstract
Missing data can significantly hamper standard time series analysis, yet in practice they are frequently encountered. In this paper, we introduce temporal Wasserstein imputation, a novel method for imputing missing data in time series. Unlike existing techniques, our approach is fully nonparametric, circumventing the need for model specification prior to imputation, making it suitable for potential nonlinear dynamics. Its principled algorithmic implementation can seamlessly handle univariate or multivariate time series with any non-systematic missing pattern. In addition, the plausible range and side information of the missing entries (such as box constraints) can easily be incorporated. As a key advantage, our method mitigates the distributional bias typical of many existing approaches, ensuring more reliable downstream statistical analysis using the imputed series. Leveraging the benign landscape of the optimization formulation, we establish the convergence of an alternating minimization algorithm to critical points. We also provide conditions under which the marginal distributions of the underlying time series can be identified. Numerical experiments, including extensive simulations covering linear and nonlinear time series models and a real-world groundwater dataset laden with missing values, corroborate the practical usefulness of the proposed method.
Keywords: Optimal transport, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
@@ 1 Introduction
In many scientific fields, data collection is often subject to various disruptions such as idiosyncratic or system-level instrument failures or changes in data collection policies, leading to missing entries in the data (Little and Rubin,, 2019). In time series analysis, missing data pose serious challenges to the application of many statistical tools since it is typically presumed that observations are available at equally spaced time points. Moreover, since the dynamic relationship is of crucial importance in time series analysis, one cannot proceed by throwing away the missing data, even if the missing data appear completely at random (i.e., the missing pattern is independent of the series).
To tackle this issue, one approach is to impute the missing entries, after which the vast array of time series methods becomes readily applicable for statistical inference in the downstream. Naturally, the estimation of the optimal interpolator, which is the expectation of the missing data conditional on the observed data, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021; McElroy and Politis,, 2020, 2022; McElroy,, 2022). Since the optimal interpolator would depend on the probabilistic structure of the underlying process, most methods involve EM-type iterations that alternate between model identification and missing data imputation (see, for example, Chpt. 4 of Peña and Tsay,, 2021). Besides the optimal interpolator, ad hoc deterministic imputation methods, such as smoothing with linear interpolation, B-splines and wavelets (see, for instance, Hastie et al.,, 2009), are also used in practice for their simplicity.
However, the optimal interpolator can cause distributional biases that are detrimental to downstream statistical analysis. To illustrate, consider an AR(1) process, $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $10<|\phi|<$ 1 and $\{\epsilon_{t}\}$ is a sequence of i.i.d. standard normal random variables. The optimal interpolator can be expressed in terms of the inverse autocorrelation functions (Peña and Maravall, 1991; McElroy and Politis, 2022; see also Chpt. 6 of Peña et al.,, 2011). For instance, suppose that data at time indices $T$ and $T+1$ are missing. Then the optimal interpolators are given by
$\hat{x}_{T}=\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T+2},\quad\hat{x}_{T+1}=\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T+2}.$
Using the fact that $\mathbb{E}(x_{t}x_{t-h})=\phi^{h}/(1-\phi^{2})$, for $h\geq 0$, it is not difficult to show that
$\mathbb{E}(\hat{x}_{T}^{2})=\mathbb{E}(\hat{x}_{T+1}^{2})=\frac{\phi^{2}+3\phi^{4}+3\phi^{6}+2\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}\mathbb{E}(x_{t}^{2})<\mathbb{E}(x_{t}^{2}),\mbox{ and}$
$|\mathbb{E}(\hat{x}_{T}\hat{x}_{T+1})|=\frac{2\phi^{2}+3\phi^{4}+3\phi^{6}+\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}|\mathbb{E}(x_{t}x_{t+1})|<|\mathbb{E}(x_{t}x_{t+1})|.$
Hence, the variance and the autocovariances of the imputed data are different from those of the original AR(1) process. In general, the joint distribution of the optimally interpolated time series is often quite different from that of the original process, leading to questionable downstream statistical analysis, especially when the number of missing data is non-negligible. Similar biases are also observed for some deterministic imputation methods (Carrizosa et al.,, 2013).
In addition to the distributional bias, the above example also reveals further predicaments of the optimal interpolator. First, in order to compute the optimal interpolator, one often needs to choose a model class before imputation. However, such a decision relies heavily on domain expertise and may be ill-informed when there are many missing values. While it is possible, for certain time series models, to estimate the model parameters without imputation, it still engages in model specification before imputation. Second, it may be difficult to compute the optimal interpolator if some nonlinear model is employed. With different missing patterns and different chosen models, the computational procedure may change drastically. Finally, incorporating side information about the missing data into this methodology appears less straightforward. Side information about the missing data restricts the set of admissible imputations. For instance, a compositional time series is known to lie in $[0,1]$ and sums to unity at each time point. Besides the optimal interpolator, many deterministic interpolators such as B-spline smoothing are also inclined to over-(or under-)shoot, leading to imputations that are outside plausible ranges.
@@
In this paper, we present the temporal Wasserstein imputation (TWI), a novel approach for time series imputation which circumvents the aforementioned limitations. As a nonparametric method, TWI is highly versatile and can be applied to a wide range of time series data. In particular, as a key advantage of TWI, it requires no model class specification prior to imputation, allowing it to perform well even when the underlying time series is nonlinear, such as the threshold autoregressive processes (TAR, Tong,, 1983). Formulated as an optimization problem, TWI can be implemented efficiently through an alternating minimization algorithm. This principled algorithmic approach allows for the seamless integration of side information by restricting the admissible imputations, and its extension to multivariate time series is also straightforward. The optimization and statistical underpinnings of TWI are closely related to the optimal transport problem (Villani,, 2008, 2021; Cuturi,, 2013), which is briefly reviewed in Section 2. Section 3 maps out the optimization formulation and introduces the computational scheme in detail.
The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series, thereby lowering the distributional bias discussed above. In Section 4, we investigate the optimization and statistical properties of TWI. Our analysis shows that the optimization landscape of TWI is biconvex and the alternating algorithm is guaranteed to converge to a critical point. However, from the statistical viewpoint, missing data imputation suffers from an identification issue: Many distributions (of the stochastic process) may be equally plausible in describing the corrupted series. This raises the question of whether and under what conditions TWI can identify the underlying marginal distributions. In Section 4.2, we show, via a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under assumptions on the missing pattern and the stability of the imputation. The latter condition motivates a simple modification of TWI, termed $k$-TWI, which shows further improvement in imputation quality from TWI in our numerical experiments.
@@
In Section 5, we evaluate the performance of TWI on synthetic data generated from a wide range of time series models, including nonlinear, multivariate, and nonstationary settings. Across all processes considered, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics such as autocovariance and model parameter estimates. Its advantage is particularly prominent in the presence of nonlinear dynamics. In Section 6, we apply TWI to a real-world groundwater dataset with many missing data. After imputation, we conduct independent component analysis and autoregressive modeling in the downstream. Compared to existing methods, TWI effectively captures the underlying dynamics, producing interpretable components as well as accurate autoregressive predictions. These numerical results support the practical contributions of TWI.
@@ 2 Preliminary: Optimal transport
[…]
If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$, for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we formally denote as $\mathcal{W}_{k}(\mu,\nu)$. Namely,
[…]
@@
The constraints in (1) enforce the joint distribution $\{\pi_{ij}\}$ to be a coupling between $\mu$ and $\nu$, and will henceforth be called the coupling constraints. In the discrete case, the optimal coupling $\{\pi_{ij}\}$ typically involves split of masses for discrete measures and is no longer the push-forward under an one-to-one map. Recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale parallelization and can be accelerated with GPUs; see Cuturi, (2013) and Peyré and Cuturi, (2019).
@@ 3 Temporal Wasserstein imputation
In this section, we introduce the proposed temporal Wasserstein imputation (TWI) method for time series imputation. First, we provide an optimization formulation of the proposed method that directly connects to the optimal transport problem. Then, we describe an alternating minimization algorithm to impute the missing values. The algorithm also reveals some connections between the proposed method and some existing approaches.
@@ 3.1 Optimization formulation
Let $\{\tilde{x}_{t}\}$ be the time series of interest and we observe the time series $\{x_{t}:t=0,1,\ldots,n-1\}$. Denote by $\bigstar$ a generic missing value, and $\mathcal{M}$ the set of time indices corresponding to the missing observations. Thus, $x_{t}=\bigstar$ if $t\in\mathcal{M}$ and $x_{t}=\tilde{x}_{t}$ otherwise. We discuss below the case that $x_{t}$ is stationary. The case of unit-root nonstationary series is discussed at the end of this section. If $\{\tilde{x}_{t}\}$ is stationary, then its $p$-dimensional marginal distribution $\mu_{p}$, i.e. the distributions of $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$, does not depend on $t$, where $p\in\mathbb{N}$. In view of this property, when imputing the missing values in $(x_{0},\ldots,x_{n-1})$, we seek an imputation $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n-1})^{\top}$ whose marginal distribution remains nearly unchanged. More specifically, we first split the time series into two parts $\{w_{0},w_{1},\ldots,w_{n_{1}}\}$ and $\{w_{n_{1}+1},w_{n_{1}+2},\ldots,w_{n-1}\}$ and define the two empirical marginal distributions
@@
$\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij},\pi_{ij}\geq 0}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t. }&\sum_{i=p-1}^{n_{1}}\pi_{ij}=\frac{1}{n-n_{1}-1}\quad\mbox{for all }j\\ &\sum_{j=n_{1}+1}^{n-1}\pi_{ij}=\frac{1}{n_{1}-p+2}\quad\mbox{for all }i,\end{aligned}\right\}$ (2)
@@
where the constraint set $\mathcal{C}$ delineates the admissible imputations. Typically, it enforces the imputation to agree with the observed data, in which case $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$. It may also reflect the side information available. For example, if $x_{t}$ are monthly sales data, the knowledge of quarterly sales define a set of linear constraints for the missing values. Another example is when $\{\mathbf{x}_{t}\}$ is a $d$-dimensional compositional time series such as the market share of several companies in an industry. It follows that $\mathbf{x}_{t}^{\top}\mathbf{1}=1$ for all $t$. Then, by defining $\mathcal{C}=\{\mathbf{W}\in\mathbb{R}^{n\times d}:\mathbf{W}\mathbf{1}_{d}=\mathbf{1}_{n}\}$, where $\mathbf{1}_{n}=(1,\ldots,1)^{\top}\in\mathbb{R}^{n}$, the set of admissible imputations are restricted to be compositional as well.
@@
where the regularization parameter $\lambda>0$ is small. As we will see in the next subsection, the $\ell_{2}$-regularization term makes a subproblem in the alternating minimization algorithm strongly convex, enhancing numerical stability. In addition, in the important special case corresponding to the Wasserstein distance of order 2, i.e., $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{2}$, the regularization allows one to solve the subproblem by finding a fixed point of a linear system. Since $L$ is often chosen to be some power of the Euclidean norm, which induces the Wasserstein distances between the marginal distributions of a time series, the solution $\hat{\mathbf{w}}$ to (3) or (4) is called the temporal Wasserstein imputation (TWI) throughout this paper.

### v3

# Temporal Wasserstein Imputation: A Versatile Method for Time Series Imputation

@@ Abstract
Missing data can significantly hamper standard time series analysis, yet they occur frequently in applications. In this paper, we introduce temporal Wasserstein imputation, a novel method for imputing missing data in time series. Unlike most existing techniques, our approach is fully nonparametric, circumventing the need for model specification prior to imputation, making it suitable for empirical applications even with nonlinear dynamics. Its principled algorithmic implementation can seamlessly handle univariate or multivariate time series with any non-systematic missing pattern. In addition, the plausible range and side information of the missing entries (such as box constraints) can easily be incorporated. Furthermore, our method mitigates the distributional bias common among many existing approaches, ensuring more reliable downstream statistical analysis using the imputed series. We establish the convergence of an alternating minimization algorithm to critical points. We also provide conditions under which the marginal distributions of the underlying time series can be identified. Numerical experiments, including extensive simulations covering both linear and nonlinear time series and an analysis on a real-world groundwater dataset, corroborate the practical usefulness of the proposed method.
Keywords: Optimal transport, Missing values, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
@@ 1 Introduction
Data collection is often subject to various disruptions at either the idiosyncratic or system level, leading to missing values. In time series analysis, missing data pose serious challenges to the application of many statistical tools because they typically assume that observations were taken at equally spaced time points. Indeed, the problem of missing data in time series has received wide attention from numerous fields, including environmental and climate studies (Afrifa-Yamoah et al.,, 2020; Junger and Ponce de Leon,, 2015), geophysics (Shen et al.,, 2015; Schoellhamer,, 2001), and engineering (Li et al.,, 2019; Racault et al.,, 2014).
Existing methods to cope with missing data abound, which can be classified into two general categories. The methods of the first category estimate the missing data before any further analysis. For example, the optimal interpolator, which is the expectation of the missing data conditional on the observed data and a given model, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021; McElroy and Politis,, 2020, 2022; McElroy,, 2022). Without any model assumption, some nonparametric imputation methods are also available. For instance, the singular spectrum analysis (SSA, Kondrashov and Ghil,, 2006; Shen et al.,, 2015; Golyandina et al.,, 2001), which relies on the spectral analysis of the trajectory matrix, is a nonparametric method to impute missing time series data. Other nonparametric methods include smoothing with polynomial interpolation or B-splines and moment matching (see Carrizosa et al.,, 2013 and Hastie et al.,, 2009). The methods of the second category circumvent the imputation stage and directly learns the model parameters. For example, by specifying a Gaussian likelihood function, one can use the Kalman filter to carry out estimation. See Harvey, (1990) and Little and Rubin, (2019), among others.
However, the aforementioned methods have some drawbacks. On the one hand, the required model specification of the optimal interpolator and the likelihood-based approaches may be ill-informed when there are many missing values. Furthermore, since computing the optimal interpolator under a nonlinear model can be challenging either computationally or analytically, simple linear models are often used in practice. Finally, the imputed series generally has a different distribution compared to the original process, which biases downstream statistical analysis especially when the number of missing data is non-negligible. On the other hand, the nonparametric SSA methods work well with structural time series featuring deterministic trends and seasonality, but are not effective in handling processes described by stochastic models, such as the celebrated ARIMA process. This is particularly so when the percentage of missing values is high.
@@
In this paper, we present the temporal Wasserstein imputation (TWI), a novel approach for time series imputation which addresses the aforementioned limitations. As a nonparametric method, TWI requires no model specification prior to imputation, allowing the underlying time series to be nonlinear, such as the threshold autoregressive process (TAR, Tong,, 1983). Hence, TWI is highly versatile and can be applied to a wide range of stochastic time series, which nicely complements the SSA-based methods. Formulated as an optimization problem, TWI can be implemented efficiently through an alternating minimization algorithm. This principled algorithmic approach not only is computationally efficient, but can also be easily generalized to handle multivariate or non-stationary time series (see Section 3). Additionally, it can seamlessly integrate the side information by restricting the admissible imputations, which can be used to enforce natural constraints such as nonnegativity constraints. The optimization and statistical underpinnings of TWI are rooted in the well-known optimal transport problem (Villani,, 2008, 2021; Cuturi,, 2013). The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series, which facilitates downstream and exploratory analyses.
Recently, optimal transport techniques have gained traction in imputation problems under different contexts. In their pioneering work, Muzellec et al., (2020) first advocated that “ideal” imputations should preserve the distribution of the underlying data. Zhao et al., (2023) called this idea “distribution matching,” and devised a deep invertible function approach to imputation, leveraging the statistical and computational properties of optimal transport (see also Wu et al.,, 2023). Hur and Liang, (2024) proposed a new method for individual treatment effect imputation based on computational optimal transport. In the time series context, after the completion of this paper we became aware of a concurrent and independent work by Wang et al., (2025), who incorporated a new Fourier-based discrepancy measure into the framework of Muzellec et al., (2020). However, because of batch sampling and the non-convexity of the optimization problem, neither Wang et al., (2025) nor Muzellec et al., (2020) has provided satisfactory optimization or statistical guarantees. In this paper, we propose an alternating minimization algorithm and study its optimization and statistical properties of TWI. Our result shows that the algorithm monotonically reduces the loss and is guaranteed to converge to a critical point. Moreover, we show, in a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under certain assumptions. This simple example encapsulates the identification problem and provides an original insight into the distribution matching framework.
@@
The rest of the paper is organized as follows. Section 2 provides a brief review of the elements of optimal transport. Section 3 maps out the optimization formulation and introduces the computational scheme of TWI in detail. It also discusses some extensions and comparisons with certain existing methods. In Section 4, we investigate the optimization and statistical properties of TWI. We also propose a simple modification of TWI based on our theoretical results, termed $k$-TWI, which shows further improvement over TWI in our numerical experiments. In Section 5, we evaluate the performance of TWI on synthetic data generated from a wide range of time series models, including nonlinear, multivariate, and nonstationary settings. Across all processes considered, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics. Its advantage is particularly prominent in the presence of nonlinear dynamics. In Section 6, we apply TWI to a real-world groundwater dataset. We use the imputed series to compute summary statistics and perform autoregressive modeling. Compared to existing methods, TWI can effectively capture the underlying dynamics, especially when the missing ratio is high.
@@ 2 Preliminary: Optimal transport
$\mathsf{OTC}(\mu,\nu)=\inf_{\pi\in\mathcal{M}(\mu,\nu)}\int L(\mathbf{x},\mathbf{y})d\pi(\mathbf{x},\mathbf{y}),$
where $\mathcal{M}$ is the set of couplings between $\mu$ and $\nu$, i.e., probability measures over $\mathbb{R}^{p}\times\mathbb{R}^{p}$ with marginals $\mu$ and $\nu$. The classic interpretation for the minimization problem is finding the most cost-effective way to move a pile of earth distributed as $\mu$ (with volume normalized to one) to fill a pit whose shape is distributed as $\nu$. It also has a matching or assignment interpretation that has led to fruitful results in the economics literature (Galichon,, 2016).
If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$, for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we denote as
$\mathcal{W}_{k}(\mu,\nu):=\left\{\inf_{\pi\in\mathcal{M}(\mu,\nu)}\int\|\mathbf{x}-\mathbf{y}\|^{k}d\pi(\mathbf{x},\mathbf{y})\right\}^{1/k}=\{\mathsf{OTC}(\mu,\nu)\}^{1/k}.$
@@
The constraints in (1) enforce the joint distribution $\{\pi_{ij}\}$ to be a coupling between $\mu$ and $\nu$, and will henceforth be called the coupling constraints. Recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale or GPU parallelization; see Cuturi, (2013) and Peyré and Cuturi, (2019).
@@ 3 Temporal Wasserstein imputation
This section introduces the temporal Wasserstein imputation (TWI) method for time series imputation. We first present its optimization formulation, which shows close connection to optimal transport. Then, we describe an alternating minimization algorithm, and highlight its relation to some existing approaches.
@@ 3.1 Optimization formulation
Let $\{\tilde{x}_{t}\}$ be the time series of interest and we observe the data $\{x_{t}:t=0,1,\ldots,n-1\}$. Denote by $\bigstar$ a generic missing value and $\mathcal{M}$ the set of time indices corresponding to the missing observations. Thus, $x_{t}=\bigstar$ if $t\in\mathcal{M}$ and $x_{t}=\tilde{x}_{t}$ otherwise. We discuss below the case that $x_{t}$ is stationary. The case of unit-root nonstationary series is discussed at the end of the next subsection. If $\{\tilde{x}_{t}\}$ is stationary, then its $p$-dimensional marginal distribution $\mu_{p}$, i.e. the distributions of $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$, does not depend on $t$, where $p\in\mathbb{N}$. In view of this property, when imputing the missing values in $(x_{0},\ldots,x_{n-1})$, we seek an imputation $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n-1})^{\top}$ whose marginal distribution remains nearly unchanged. More specifically, we first split the time series into two parts $\{w_{0},w_{1},\ldots,w_{n_{1}}\}$ and $\{w_{n_{1}+1},w_{n_{1}+2},\ldots,w_{n-1}\}$ and define the two empirical marginal distributions
@@
$\left.\begin{aligned} \mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij},\pi_{ij}\geq 0}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t. }&\sum_{i=p-1}^{n_{1}}\pi_{ij}=\frac{1}{n-n_{1}-1}\quad\mbox{for all }j\\ &\sum_{j=n_{1}+1}^{n-1}\pi_{ij}=\frac{1}{n_{1}-p+2}\quad\mbox{for all }i,\end{aligned}\right\}$ (2)
@@
where the constraint set $\mathcal{C}$ delineates the admissible imputations. Typically, it ensures the imputation to agree with the observed data, in which case $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$. It may also reflect the side information available. For example, if $x_{t}$ are monthly sales data, the knowledge of quarterly sales define a set of linear constraints for the missing values. Another example is when $\{\mathbf{x}_{t}\}$ is a $d$-dimensional compositional time series such as the market share of several companies in an industry. It follows that $\mathbf{x}_{t}^{\top}\mathbf{1}=1$ for all $t$. Then, by defining $\mathcal{C}=\{\mathbf{W}\in\mathbb{R}^{n\times d}:\mathbf{W}\mathbf{1}_{d}=\mathbf{1}_{n}\}$, where $\mathbf{1}_{n}=(1,\ldots,1)^{\top}\in\mathbb{R}^{n}$, the set of admissible imputations are restricted to be compositional as well.
@@
where the regularization parameter $\lambda>0$ is small. As we will see in the next subsection, the $\ell_{2}$-regularization term enhances the numerical stability of the alternating minimization algorithm. Since $L$ is often chosen to induce the Wasserstein distances between the marginal distributions of a time series, the solution $\hat{\mathbf{w}}$ to (3) or (4) is called the temporal Wasserstein imputation (TWI) throughout this paper.